OpenStudy (ledah):

What is the lateral area of a square pyramid whose base side is 7.5 ft. and slant height is 16 ft?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {p \times l \over2 }}\)

Parth (parthkohli):

p = Perimeter of the square. l = Slant height.

OpenStudy (ledah):

so would I add 7.5, four times and then multiply by 16?

Parth (parthkohli):

Yep. Good. And then you'd divide the number you get by 2.

OpenStudy (ledah):

150

Parth (parthkohli):

How so?

OpenStudy (ledah):

I got 300 from adding the numbers then I divided which gave me 150

Parth (parthkohli):

Well. \(7.5 \times 4 = P = 30\)

Parth (parthkohli):

What is \(30 \times 16\)?

OpenStudy (ledah):

Oh

OpenStudy (ledah):

480

Parth (parthkohli):

Yep. 480/2 = ?

OpenStudy (ledah):

240

Parth (parthkohli):

That's it :)

OpenStudy (ledah):

thanks parth

Parth (parthkohli):

No problem.